find the approximate value of ∛127


1
Expert's answer
2022-01-17T13:05:31-0500

Let y=x1/3

Let x=125 and "x+\Delta x=127"

"\Delta x=2"

y=1251/3=5

"\frac{dy}{dx}=1/3(x)^{-2/3}"

"\Delta y=\frac{x^{-2/3}}{3}\Delta x=\frac{2}={3(5)^2}=0.0266"

"y+\Delta y=5+0.0266=5.0266"